OpenStudy (anonymous):

Write the equation of a line with slope m = 3 and including point (3, 5).

OpenStudy (zzr0ck3r):

y = mx + b y = 3x+b 5 = 3*3+b 5 = 9 +b b = -4 y = 3x-4

OpenStudy (zzr0ck3r):

make since?

OpenStudy (anonymous):

yep got it thanks